OpenStudy (141823):

A bag of nickels and dimes contain 83 coins. If the value of the coins was $7.30, how many nickels are there? (Hint: If you call the number of nickels x, then the value of the nickels is 5x. If you call the number of dimes y, then the value of the dimes is 10y) Medal will be given!

OpenStudy (anonymous):

0.5x + .10y = 7.30 there are 83 coins, that is x + y = 83 solving this equation for x we have x = 83 - y plug that in for x in your original equation 0.5(83-y) + .10y = 7.30 solve for y

OpenStudy (anonymous):

once you solve for y, then you can plug that into x + y = 83 and solve for x